# 2cm2cm1cm1cm

### v4

@@ 2.3 Diagonalization of Quadratic Hamiltonians
with $D_{0}\left(\Upsilon_{0}^{\top}\right)^{-1}D_{0}^{\ast}$ being a well-defined operator acting on $\mathfrak{h}$. Up to singular cases, these assumptions may be necessary for the convergence of the flow because the existence of $\Upsilon_{\infty}$ seems to generally imply its positivity (up to singular cases). See the spectral properties of the operator family (195). They correspond to take $\varepsilon,\mu\rightarrow 0$ in the conditions of Theorem 2.4 with a possible limit operator $\Upsilon_{\infty}\doteq\Upsilon_{0}+\Delta_{\infty}>0$ having a priori no spectral gap. The conditions expressed by (18) look similar to the situation $\Omega_{0}>0$, without spectral gap, studied in the hyperbolic case [6]. By analogy with the hyperbolic case, one may assume that $\left(\Upsilon_{0}\right)^{-1/2}D_{0}$ is a Hilbert-Schmidt operator to ensure the existence of $\Delta_{\infty}$ as a trace-class operator. The arguments to handle the non-spectral gap case in [6] were technically involved and long (with domain issues) and the analogous study in the fermionic case under Condition (18) has probably the same mathematical complexity.
@@ 2.4 Araki’s Approach on the Level of $C^{\ast}$-Algebras
In 1968, Araki established a very general method for some “$\mathrm{N}$–diagonalization” of bilinear Hamiltonians. This refers to the well-known paper [18]. Such a diagonalization holds true in great generality. The Hilbert-Schmidt condition $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$ and the semiboundedness of $\Upsilon_{0}$ are not even required in [18, Theorem 5.4]. This is because the concepts studied in Araki’s paper [18] are different than those analyzed here, albeit still related. Araki’s terminology and notation used in [18] are almost identical to the ones used in other works like in Berezin’s book [11], but they refer to differents concepts or frameworks. For instance, Araki’s study [18] uses a purely $C^{\ast}$-algebraic formulation, while Berezin and we work on the Fock space representation. See also [30] for more details on Araki’s works in relation with the Fock space representation.
@@
[…]
@@
  • A Bogoliubov transformation is a unitary operator $U\in\mathcal{B}(\mathcal{H})$ such that $U\mathfrak{A=A}U$. A self-dual Hamiltonian $h$ on $\mathcal{H}$ is a self-adjoint operator satisfying $h=-\mathfrak{A}h\mathfrak{A}$ with domain $\mathcal{D}(h)=\mathfrak{A}(\mathcal{D}(h))\subseteq\mathcal{\mathcal{H}}$. Basis projections are orthogonal projections $P\in\mathcal{B}(\mathcal{H})$ satisfying $\mathfrak{A}P\mathfrak{A}=P^{\bot}\doteq\mathbf{1}_{\mathcal{H}}-P$.
  • The Bogoliubov $\ast$–automorphism associated with a Bogoliubov transformation $U\in\mathcal{B}(\mathcal{H})$ is the unique $\ast$–automorphism $\mathbf{\chi}_{U}$ of $\mathcal{A}$ satisfying
    $\mathbf{\chi}_{U}\left(\mathrm{B}(\varphi)\right)=\mathrm{B}(U\varphi)\ ,\qquad\varphi\in\mathcal{\mathcal{H}}\ .$ (23)
@@
    keeping in mind that [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] usually denotes a commutator. Bilinear Hamiltonians are by definition such a derivation, or such a “$\mathrm{H}$”. The bilinear Hamiltonian “$\mathrm{H}$” is said hermitian or self-adjoint because $h=h^{\ast}$. See [18, Definition 5.1].
@@
The motivation of these definitions comes from the fact that such a $\mathrm{H}$ in Araki’s viewpoint can be seen in some very restricted situation as a quadratic Hamiltonian in our sense. See, e.g., [18, Remark 4.3] and [33, Sections 2.2-2.3], where $\mathrm{H}$ is shown to exist as a self-adjoint element of the $C^{\ast}$-algebra $\mathcal{A}$ for trace-class Hamiltonians $h$. However, in general, “$\mathrm{H}$” is only a formal notation, without necessarily another meaning than denoting the symmetric derivation via (24). In other words, [18] analyzes another situation and, in the general case, it is an open problem to show the existence of some well-defined, self-adjoint, quadratic Hamiltonian that leads to $\delta_{h}$ in some representation. In the Fock space representation, one can answer this question provided (4) defines a self-adjoint operator on the Fock space:
@@
for any $\psi\in\mathcal{H}=\mathfrak{h}\oplus\mathfrak{h}$. The proof is relatively standard in its principle. See for instance [33, Lemma 2.8]. We outline it for completeness. As usual, $[A,B]\doteq AB-BA$ is the commutator of two elements $A,B\in\mathcal{A}$. Using the CAR (6) (or (20)), Equation (22) together with the properties of the antiunitary involution $\mathfrak{A}$, $h^{\ast}=-\mathfrak{A}h\mathfrak{A}$ and $\mathrm{B}(\psi)^{\ast}=\mathrm{B}(\mathfrak{A}(\psi))$, we compute that
@@
[18, Theorem 5.4] would mean in this case that a quadratic Hamiltonian in the sense of (4) is diagonalizable whenever (25) is essentially self-adjoint and such that the eigenspace associated with the eigenvalue 00 has dimension even or infinite. This would be a very general result, even if $D_{0}$ is required to be Hilbert-Schmidt to ensure the assertions of Lemma 2.7. For instance, the operator $\Upsilon_{0}$ is not even semibounded in this case. However, [18, Definition 5.1], defining the diagonalization of bilinear Hamiltonians, only involves the existence of a Bogoliubov transformation $U\in\mathcal{B}(\mathcal{H})$ such that $U\mathfrak{A=A}U$, which may not define a unitary transformation in the Fock space representation. This is well known via the celebrated Shale-Stinespring condition which refers to the Hilbert-Schmidt assumption.
@@
Let $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ and $\mathfrak{A}$ be defined by (21). Take any Bogoliubov transformation $U$. In particular, $U$ has the form [^9] Use $\overline{X}=\mathcal{C}X\mathcal{C}$ (Remark 2.1) and $U\mathfrak{=A}U\mathfrak{A}$ together with (21). The fact that $UU^{\ast}=U^{\ast}U=1$ also implies other constraints on $\mathbf{u},\mathbf{v}\in\mathcal{B}\left(\mathfrak{h}\right)$, but this is not important here.
$U=\left(\begin{array}[]{cc}\mathbf{u}&\mathbf{v}\\ \overline{\mathbf{v}}&\overline{\mathbf{u}}\end{array}\right)$
[…]
@@
See [16, Theorems 3.1 and 6.2] or [17, Theorem 7] for a complete proof. This theorem originally goes back to Shale and Stinespring [34]. For more recent or pedagogical references, see [35, Theorem 9.5] or [15, Equation (2a.9) and Theorem 2.2]. Again the Hilbert-Schmidt condition is referred to as the Shale-Stinespring condition. When it is satisfied, one says that $U$ is (unitarily) implementable. Note that one can extend this result to the case where $\mathrm{U}$ is not necessarily unitary. This refers to [36, Proposition 1.2].
@@ 2.5 Brockett-Wegner Flow and Elliptic Operator Flows
The $\mathrm{N}$–diagonalization of fermionic quadratic Hamiltonians may be shown by using an approach that could be similar to [8], done in the bosonic case. However, we follow here another path by using the Brockett-Wegner flow [19, 20, 21, 22] as a guideline. This method of proof is interesting in its own right. Indeed, the Brockett-Wegner flow is very little developed in Mathematics, although it has been often used in theoretical physics [23]. In our setting, it allows one to connect the $\mathrm{N}$–diagonalization of fermionic quadratic Hamiltonians to an elegant non-linear elliptic flow on operators acting on the one-particle Hilbert space, which is studied in detail in [24] (see Section A).
In this section, we thus formally explain this connection. This allows the interested reader to understand the general strategy to prove Theorem 2.4, before going through the rigorous arguments performed in Section 3. In addition, this demonstrates how the formal Brockett-Wegner flow can produce novel and elegant ordinary differential equations (ODEs) on operator spaces. Remark in this context that differential equations on non-commutative spaces (e.g., operator spaces) are clearly much less known than its commutative counterpart (i.e., ODEs or PDEs on function spaces), which forms a well-established, albeit still very active, research field of Mathematics. This section can thus be viewed as an invitation to develop this research direction on operator spaces.
@@
with $[A,B]\doteq AB-BA$ being the commutator between two operators $A$ and $B$ (provided it exists on some domain). This operator flow is very useful because it is awaited that the solution to (30) is a family of unitarily equivalent quadratic operators: For any $s\in\mathbb{R}_{0}^{+}$ and $t\in\left[s,\infty\right)$, $\mathrm{H}_{t}=\mathrm{U}_{t,s}\mathrm{\mathrm{H}}_{s}\mathrm{U}_{t,s}^{\ast}$ with $\left(\mathrm{U}_{t,s}\right)_{t\geq s}$ being a (evolution) family of unitary operators solution to the non-autonomous (hyperbolic in Kato’s terminology [28, 29], named Kato-hyperbolic here to avoid confusions) evolution equation
@@
$\forall s\in\mathbb{R}_{0}^{+},\ t\in\left[s,\infty\right):\qquad\partial_{t}U_{t,s}=-ig_{t}U_{t,s}\ ,\qquad U_{s,s}\doteq\mathbf{1}\ .$ (34)
By (21) and $D_{t}=-D_{t}^{\top}$ together with Remark 2.1, $g_{t}\mathfrak{A}=\mathfrak{A}g_{t}$ for all $t\in\mathbb{R}_{0}^{+}$ and, via the uniqueness of the solution to the above evolution equation, one can check that $\left(U_{t,s}\right)_{t\geq s}$ is a family of Bogoliubov transformations. In particular, $U_{t,s}\mathfrak{A=A}U_{t,s}$. Moreover, assuming a unique solution to the flow (32), we should have $h_{t}=U_{t,0}h_{0}U_{t,0}^{\ast}$ for all $t\in\mathbb{R}_{0}^{+}$. For instance, the Hilbert-Schmidt norm of $h_{(\cdot)}$ (if it exists) is constant, which can be rewritten as
@@
The Brockett-Wegner flow (32) takes place in the one-particle space, but we have to lift it again to the Fock space. In particular, the above family $\left(U_{t,s}\right)_{t\geq s}$ of Bogoliubov transformations has to be implemented, as explained in Remark 2.9. Mutatis mutandis for $\left(h_{t}\right)_{t\geq 0}$ and the relation $h_{t}=U_{t,0}h_{0}U_{t,0}^{\ast}$. This is done in Sections 3.2 and 3.3 whereas we present results on general implementation criteria of quadratic Hamiltonians in Section 2.6.
@@ 2.6 Bach, Lieb and Solovej’s Approach to Quadratic Hamiltonians
Bach, Lieb and Solovej [15] introduced a definition of quadratic Hamiltonians that a priori differs from the one used here. This definition is inspired by Araki’s definition [18, Definition 5.1] and Theorem 2.8. For the reader’s convenience, it is reproduced below within the terminology and notation used in Section 2.4. We take here $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ and a Bogoliubov transformation is a unitary operator $U\in\mathcal{B}(\mathcal{H})$ such that $U\mathfrak{A=A}U$.
@@
A self-adjoint operator $\mathrm{H}$ on the fermionic Fock space $\mathcal{F}$ is said to be a quadratic Hamiltonian if the unitary operator $\mathrm{U}_{t}=\mathrm{e}^{it\mathrm{H}/2}$ implements a Bogoliubov transformation $U_{t}$ for any $t\in\mathbb{R}$, i.e.,
$\mathrm{U}_{t}\left(\mathbf{\pi}_{\mathcal{F}}\left(A\right)\right)\mathrm{U}_{t}^{\ast}=\mathbf{\pi}_{\mathcal{F}}\circ\mathbf{\chi}_{U_{t}}\left(A\right)\ ,\qquad A\in\mathcal{A},\ t\in\mathbb{R}\ ,$ (37)
where $\mathbf{\chi}_{U_{t}}$ is the unique $\ast$–automorphism $\mathbf{\chi}_{U}$ of $\mathcal{A}$ satisfying (23) for $U=U_{t}$ while $\mathbf{\pi}_{\mathcal{F}}$ is the Fock space representation of the universal $C^{\ast}$-algebra $\mathcal{A}$ (19).
This point of view is also studied for bosonic systems, for instance by Bruneau and Derezinski via their definition of Bogoliubov Hamiltonians [7]. Observe also that, if $\mathrm{H}$ is a quadratic Hamiltonian as in Definition 2.10 then, for any real constant $C\in\mathbb{R}$$\mathrm{H}+C\mathbf{1}$ is of course a quadratic Hamiltonian leading to the same Bogoliubov transformation.
@@
Fix $E_{0}\in\mathbb{R}$. Let $\Upsilon_{0}=\Upsilon_{0}^{\ast}$ and $D_{0}=-D_{0}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ be two operators acting on the Hilbert space $\mathfrak{h}$. Then, the self-adjoint extension $\mathrm{H}_{0}\equiv\mathrm{H}_{0}^{\ast\ast}$ of the quadratic operator defined by (4)–(5) on the domain (5) is a quadratic Hamiltonian with respect to Definition 2.10.
@@
We now discuss whether any quadratic Hamiltonian with respect to Definition 2.10 is the self-adjoint extension of a quadratic operators defined by (4) on the domain (5). To this end, we recall that $\mathfrak{A}$ is the antiunitary involution of $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ defined by (21), while $\{\mathrm{B}(\psi)\}_{\psi\in\mathcal{\mathcal{H}}}$ is the family of generators of the universal $C^{\ast}$-algebra $\mathcal{A}$ (19). We start with a first elementary, but important, lemma:

### v6

@@ 2.3 Diagonalization of Quadratic Hamiltonians
with $D_{0}\left(\Upsilon_{0}^{\top}\right)^{-1}D_{0}^{\ast}$ being a well-defined operator acting on $\mathfrak{h}$. Up to singular cases, these assumptions may be necessary for the convergence of the flow because the existence of $\Upsilon_{\infty}$ seems to generally imply its positivity (up to singular cases). See the spectral properties of the operator family studied in [24, Theorem 3]. They correspond to take $\varepsilon,\mu\rightarrow 0$ in the conditions of Theorem 2.4 with a possible limit operator $\Upsilon_{\infty}\doteq\Upsilon_{0}+\Delta_{\infty}\geq 0$ having a priori no spectral gap. The conditions expressed by (18) look similar to the situation $\Omega_{0}\geq 0$, without spectral gap, studied in the hyperbolic case [6]. By analogy with the hyperbolic case, one may assume that $\left(\Upsilon_{0}\right)^{-1/2}D_{0}$ is a Hilbert-Schmidt operator to ensure the existence of $\Delta_{\infty}$ as a trace-class operator. The arguments to handle the non-spectral gap case in [6] were technically involved and long (with domain issues). We do not expect any fundamental obstruction for the analogous study in the fermionic case under Condition (18), but we refrain from doing so, limiting in particular the length of this document.
@@ 2.4 Araki’s Approach on the Level of $C^{\ast}$-Algebras
In 1968, Araki established a very general method for some $\mathrm{N}$–diagonalization” of bilinear Hamiltonians. This refers to the well-known paper [18]. Such a diagonalization holds true in great generality. The Hilbert-Schmidt condition $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$ and the semiboundedness of $\Upsilon_{0}$ are not even required in [18, Theorem 5.4]. This is because the concepts studied in Araki’s paper [18] are different than those analyzed here, albeit still related. Araki’s terminology and notation used in [18] are almost identical to the ones used in other works like in Berezin’s book [11], but they refer to differents concepts or frameworks. For instance, Araki’s study [18] uses a purely $C^{\ast}$-algebraic formulation, while Berezin and we work on the Fock space representation. See also [30] for more details on Araki’s works in relation with the Fock space representation.
@@
    $\mathrm{B}(\psi_{1})\mathrm{B}(\psi_{2})^{\ast}+\mathrm{B}(\psi_{2})^{\ast}\mathrm{B}(\psi_{1})=\left\langle\psi_{1},\psi_{2}\right\rangle_{\mathcal{H}}\,\mathfrak{1}\ .$ (20)
@@
  • A Bogoliubov transformation is a unitary operator $\mathcal{U}\in\mathcal{B}(\mathcal{H})$ such that $\mathcal{U}\mathfrak{A=A}\mathcal{U}$. A self-dual Hamiltonian $h$ on $\mathcal{H}$ is a self-adjoint operator satisfying $h=-\mathfrak{A}h\mathfrak{A}$ with domain $\mathcal{D}(h)=\mathfrak{A}(\mathcal{D}(h))\subseteq\mathcal{\mathcal{H}}$. Basis projections are orthogonal projections $P\in\mathcal{B}(\mathcal{H})$ satisfying $\mathfrak{A}P\mathfrak{A}=P^{\bot}\doteq\mathbf{1}_{\mathcal{H}}-P$.
  • The Bogoliubov $\ast$–automorphism associated with a Bogoliubov transformation $\mathcal{U}\in\mathcal{B}(\mathcal{H})$ is the unique $\ast$–automorphism $\mathbf{\chi}_{\mathcal{U}}$ of $\mathcal{A}$ satisfying
    $\mathbf{\chi}_{\mathcal{U}}\left(\mathrm{B}(\varphi)\right)=\mathrm{B}(\mathcal{U}\varphi)\ ,\qquad\varphi\in\mathcal{\mathcal{H}}\ .$ (23)
@@
    keeping in mind that $\cdot$$\cdot$] usually denotes a commutator. Bilinear Hamiltonians are by definition such a derivation, or such a $\mathrm{H}$”. The bilinear Hamiltonian $\mathrm{H}$” is said hermitian or self-adjoint because $h=h^{\ast}$. See [18, Definition 5.1].
@@
The motivation of these definitions comes from the fact that such a $\mathrm{H}$ in Araki’s viewpoint can be seen in some very restricted situation as a quadratic Hamiltonian in our sense. See, e.g., [18, Remark 4.3] and [33, Sections 2.2-2.3], where $\mathrm{H}$ is shown to exist as a self-adjoint element of the $C^{\ast}$-algebra $\mathcal{A}$ for trace-class Hamiltonians $h$. However, in general, $\mathrm{H}$” is only a formal notation, without necessarily another meaning than denoting the symmetric derivation via (24). In other words, [18] analyzes another situation and, in the general case, it is an open problem to show the existence of some well-defined, self-adjoint, quadratic Hamiltonian that leads to $\delta_{h}$ in some representation. In the Fock space representation, one can answer this question provided (4) defines a self-adjoint operator on the Fock space:
@@
for any $\psi\in\mathcal{H}=\mathfrak{h}\oplus\mathfrak{h}$. The proof is relatively standard in its principle. See for instance [33, Lemma 2.8]. We outline it for completeness. As usual, $A,B]\doteq AB-BA$ is the commutator of two elements $A,B\in\mathcal{A}$. Using the CAR (6) (or (20)), Equation (22) together with the properties of the antiunitary involution $\mathfrak{A}$, $h^{\ast}=-\mathfrak{A}h\mathfrak{A}$ and $\mathrm{B}(\psi)^{\ast}=\mathrm{B}(\mathfrak{A}(\psi))$, we compute that
@@
[18, Theorem 5.4] would mean in this case that a quadratic Hamiltonian in the sense of (4) is diagonalizable whenever (25) is essentially self-adjoint and such that the eigenspace associated with the eigenvalue 0 has dimension even or infinite. This would be a very general result, even if $D_{0}$ is required to be Hilbert-Schmidt to ensure the assertions of Lemma 2.7. For instance, the operator $\Upsilon_{0}$ is not even semibounded in this case. However, [18, Definition 5.1], defining the diagonalization of bilinear Hamiltonians, only involves the existence of a Bogoliubov transformation $\mathcal{U}\in\mathcal{B}(\mathcal{H})$ such that $\mathcal{U}\mathfrak{A=A}\mathcal{U}$, which may not define a unitary transformation in the Fock space representation. This is well known via the celebrated Shale-Stinespring condition which refers to the Hilbert-Schmidt assumption.
@@
Let $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ and $\mathfrak{A}$ be defined by (21). Take any Bogoliubov transformation $\mathcal{U}$. In particular, $\mathcal{U}$ has the form [^11] Use $\overline{X}=\mathcal{C}X\mathcal{C}$ (Remark 2.1) and $\mathcal{U}=\mathfrak{A}\mathcal{U}\mathfrak{A}$ together with (21). The fact that $\mathcal{U}\mathcal{U}^{\ast}=\mathcal{U}^{\ast}\mathcal{U}=1$ also implies other constraints on $\mathbf{u},\mathbf{v}\in\mathcal{B}\left(\mathfrak{h}\right)$, but this is not important here.
$\mathcal{U}=\left(\begin{array}[]{cc}\mathbf{u}&\mathbf{v}\\ \overline{\mathbf{v}}&\overline{\mathbf{u}}\end{array}\right)$
with $\mathbf{u},\mathbf{v}\in\mathcal{B}\left(\mathfrak{h}\right)$. Then there is a unitary transformation $\mathrm{U}:\mathcal$$\rightarrow F$} such that $\mathrm{U}\left(\mathbf{\pi}_{\mathcal{F}}\left(A\right)\right)\mathrm{U}^{\ast}=\mathbf{\pi}_{\mathcal{F}}\circ\mathbf{\chi}_{\mathcal{U}}\left(A\right)$ for all $A\in\mathcal{A}$ iff $\mathbf{v}\in\mathcal{L}^{2}(\mathfrak{h})$. In particular, by Equation (22), when $\mathbf{v}\in\mathcal{L}^{2}(\mathfrak{h})$,
@@
See [16, Theorems 3.1 and 6.2] or [17, Theorem 7] for a complete proof. This theorem originally goes back to Shale and Stinespring [34]. For more recent or pedagogical references, see [35, Theorem 9.5] or [15, Equation (2a.9) and Theorem 2.2]. Again the Hilbert-Schmidt condition is referred to as the Shale-Stinespring condition. When it is satisfied, one says that $\mathcal{U}$ is (unitarily) implementable. Note that one can extend this result to the case where $\mathrm{U}$ is not necessarily unitary. This refers to [36, Proposition 1.2].
@@ 2.5 Brockett-Wegner Flow and Elliptic Operator Flows
We propose to tackle the diagonalization of quadratic fermionic Hamiltonians by using the Brockett-Wegner flow [19, 20, 21, 22] as a guideline. This method of proof is interesting in its own right. Indeed, the Brockett-Wegner flow is very little developed in Mathematics, although it has been often used in theoretical physics [23]. In our setting, it allows one to connect the $\mathrm{N}$–diagonalization of fermionic quadratic Hamiltonians to an elegant non-linear elliptic flow on operators acting on the one-particle Hilbert space, which is studied in detail in [24] (see Section A.1).
In this section, we thus formally explain this connection. This allows the interested reader to understand the general strategy to prove Theorem 2.4, before going through the rigorous arguments performed in Section 3. In addition, this demonstrates how the formal Brockett-Wegner flow can produce novel and elegant ordinary differential equations (ODEs) on operator spaces and this section can thus be viewed as an invitation to develop this research direction.
@@
with $A,B]\doteq AB-BA$ being the commutator between two operators $A$ and $B$ (provided it exists on some domain). This operator flow is very useful because it is awaited that the solution to (30) is a family of unitarily equivalent quadratic operators: For any $s\in\mathbb{R}_{0}^{+}$ and $t\in\left[s,\infty\right)$, $\mathrm{H}_{t}=\mathrm{U}_{t,s}\mathrm{\mathrm{H}}_{s}\mathrm{U}_{t,s}^{\ast}$ with $\left(\mathrm{U}_{t,s}\right)_{t\geq s}$ being a (evolution) family of unitary operators solution to the non-autonomous (hyperbolic in Kato’s terminology [28, 29], named Kato-hyperbolic here to avoid confusions) evolution equation
@@
$\forall s\in\mathbb{R}_{0}^{+},\ t\in\left[s,\infty\right):\qquad\partial_{t}\mathcal{U}_{t,s}=-ig_{t}\mathcal{U}_{t,s}\ ,\qquad\mathcal{U}_{s,s}\doteq\mathbf{1}\ .$ (34)
By (21) and $D_{t}=-D_{t}^{\top}$ together with Remark 2.1, $g_{t}\mathfrak{A}=\mathfrak{A}g_{t}$ for all $t\in\mathbb{R}_{0}^{+}$ and, via the uniqueness of the solution to the above evolution equation, one can check that $\left(\mathcal{U}_{t,s}\right)_{t\geq s}$ is a family of Bogoliubov transformations. In particular, $\mathcal{U}_{t,s}\mathfrak{A=A}\mathcal{U}_{t,s}$. Moreover, assuming a unique solution to the flow (32), we should have $h_{t}=\mathcal{U}_{t,0}h_{0}\mathcal{U}_{t,0}^{\ast}$ for all $t\in\mathbb{R}_{0}^{+}$. For instance, the Hilbert-Schmidt norm of $h_{(\cdot)}$ (if it exists) is constant, which can be rewritten as
@@
The Brockett-Wegner flow (32) takes place in the one-particle space, but we have to lift it again to the Fock space. In particular, the above family $\left(\mathcal{U}_{t,s}\right)_{t\geq s}$ of Bogoliubov transformations has to be implemented, as explained in Remark 2.9. Mutatis mutandis for $\left(h_{t}\right)_{t\geq 0}$ and the relation $h_{t}=\mathcal{U}_{t,0}h_{0}\mathcal{U}_{t,0}^{\ast}$. This is done in Sections 3.2 and 3.3 whereas we present results on general implementation criteria of quadratic Hamiltonians in Section 2.6.
@@ 2.6 Bach, Lieb and Solovej’s Approach to Quadratic Hamiltonians
Bach, Lieb and Solovej [15] introduced a definition of quadratic Hamiltonians that a priori differs from the one used here. This definition is inspired by Araki’s definition [18, Definition 5.1] and Theorem 2.8. For the reader’s convenience, it is reproduced below within the terminology and notation used in Section 2.4. We take here $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ and a Bogoliubov transformation is a unitary operator $\mathcal{U}\in\mathcal{B}(\mathcal{H})$ such that $\mathcal{U}\mathfrak{A=A}\mathcal{U}$.
@@
A self-adjoint operator $\mathrm{H}$ on the fermionic Fock space $\mathcal{F}$ is said to be a quadratic Hamiltonian if the unitary operator $\mathrm{U}_{t}=\mathrm{e}^{it\mathrm{H}/2}$ implements a Bogoliubov transformation $\mathcal{U}_{t}$ for any $t\in\mathbb{R}$, i.e.,
$\mathrm{U}_{t}\left(\mathbf{\pi}_{\mathcal{F}}\left(A\right)\right)\mathrm{U}_{t}^{\ast}=\mathbf{\pi}_{\mathcal{F}}\circ\mathbf{\chi}_{\mathcal{U}_{t}}\left(A\right)\ ,\qquad A\in\mathcal{A},\ t\in\mathbb{R}\ ,$ (37)
where $\mathbf{\chi}_{\mathcal{U}_{t}}$ is the unique $\ast$–automorphism $\mathbf{\chi}_{\mathcal{U}}$ of $\mathcal{A}$ satisfying (23) for $\mathcal{U}=\mathcal{U}_{t}$ while $\mathbf{\pi}_{\mathcal{F}}$ is the Fock space representation of the universal $C^{\ast}$-algebra $\mathcal{A}$ (19).
This point of view is also studied for bosonic systems, for instance by Bruneau and Derezinski via their definition of Bogoliubov Hamiltonians [7]. Observe also that, if $\mathrm{H}$ is a quadratic Hamiltonian as in Definition 2.10 then, for any real constant $C\in\mathbb{R}$, $\mathrm{H}+C\mathbf{1}$ is of course a quadratic Hamiltonian leading to the same Bogoliubov transformation.
@@
Fix $E_{0}\in\mathbb{R}$. Let $\Upsilon_{0}=\Upsilon_{0}^{\ast}$ and $D_{0}=-D_{0}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ be two operators acting on the Hilbert space $\mathfrak{h}$. Then, the self-adjoint extension $\mathrm{H}_{0}\equiv\mathrm{H}_{0}^{\ast\ast}$ of the quadratic operator defined by (4)–(5) on the domain (5) is a quadratic Hamiltonian in the sense of Definition 2.10.
@@
We now discuss whether any quadratic Hamiltonian in the sense of Definition 2.10 is the self-adjoint extension of a quadratic operators defined by (4) on the domain (5). To this end, we recall that $\mathfrak{A}$ is the antiunitary involution of $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ defined by (21), while $\{\mathrm{B}(\psi)\}_{\psi\in\mathcal{\mathcal{H}}}$ is the family of generators of the universal $C^{\ast}$-algebra $\mathcal{A}$ (19). We start with a first elementary, but important, lemma: